# Flat quasi-Frobenius Lie superalgebras

### v1

@@ 1.1. The Riemannian setting
The curvature of $(\mathfrak{g},\langle\;,\;\rangle)$ at the identity is
@@
A pseudo-Riemannian Lie group $(G,\mu)$ is called flat when its curvature vanishes identically; in this case, $(\mathfrak{g},\langle\;,\;\rangle)$ will be refered to as a flat pseudo-Riemannian Lie algebra. In fact, flatness is equivalent to $(\mathfrak{g},\star)$ being a left-symmetric algebra..
@@
As in the pseudo-Riemannian case, one defines the curvature of this connection. A symplectic Lie group $(G,\Omega)$ is called flat when its curvature vanishes identically. In this case, the induced Lie algebra $(\mathfrak{g},\omega)$ is a flat quasi-Frobenius Lie algebra. In fact, flatness is equivalent to the fact that the symplectic product defines a left-symmetric structure on $\mathfrak{g}$.
@@ 1.3. Outline of the paper
for all homogeneous $u,v,w\in{\mathcal{A}}$. The product $\bullet$ is also characterized by a Koszul formula (see Eq. 4). Its existence and uniqueness are established in Prop 3.1.1.  Proposition.
Further, we study the Levi-Civita product in the case where the pre-symplectic superalgebra is a Lie superalgebra and the form $\omega$ is closed. We call these superalgebras quasi-Frobenius Lie superalgebras, following [BBE, BE, BM]. It is worth mentioning that several authors call them symplectic Lie superalgebras (see [BC, BEB, BBNS] to cite a few). We show that the Levi-Civita product associated with a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is always right-symmetric (see Prop. 3.1.5.  Proposition) and satisfies
@@
In addition, we introduce a second product $\star$ on a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$, defined by
@@
We show nevertheless that in any quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ we can construct a unique symplectic product with the key property that it depends only on the Lie bracket and the symplectic form $\omega$. This product is explicitly given by:
@@
In Section 6, we show that all non-abelian flat quasi-Frobenius Lie superalgebras have a degenerate center, see Theorem 6.1.  Theorem. Moreover, every such Lie superalgebra can be obtained via flat double extensions. In particular, every flat orthosymplectic quasi-Frobenius Lie superalgebra can be constructed by a sequence of flat double extensions starting from an an abelian Lie superalgebra (see Corollary 6.8.  Corollary), while flat periplectic quasi-Frobenius Lie superalgebras can be obtained by a sequence of flat double extensions starting from the trivial algebra $\{0\}$ (see Corollary 6.9.2.  Corollary). Furthermore, we show that every flat orthosymplectic quasi-Frobenius Lie superalgebra is not only solvable but nilpotent (Theorem 6.9.  Theorem).
@@ 2. Backgrounds
$\omega(u,v)=-(-1)^{|u||v|}\omega(v,u)\text{ for every }u,v\in V.$
@@
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a finite-dimensional superspace with $V_{\bar{1}}\neq\{0\}$, endowed with a periplectic form $\omega$. Then, $\dim(V_{\bar{0}})=\dim(V_{\bar{1}})$.
@@
$N_{\ell}(\mathfrak{g},\bullet):=\left\{u\in\mathfrak{g},\mathrm{L}^{\bullet}_{u}=0\right\},\;N_{r}(\mathfrak{g},\bullet):=\left\{u\in\mathfrak{g},\mathrm{R}_{u}^{\bullet}=0\right\},\;\text{and}\;N(\mathfrak{g},\bullet):=N_{\ell}(\mathfrak{g},\bullet)\cap N_{r}(\mathfrak{g},\bullet).$
@@ 3.1. The Levi-Civita product
(Uniqueness) Suppose that there exist two products $\bullet$ and $\bullet^{\prime}$ satisfying conditions (2), (3). For all $u,v,w\in\mathfrak{g}$, we have
$\omega(u\bullet v,w)$ $=\omega([u,v],w)+(-1)^{|u||v|}\omega(v\bullet u,w)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega(v\bullet w,u)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)-(-1)^{|u||v|+|w||u|+|v||w|}\omega(w\bullet v,u)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega(w\bullet u,v)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega([w,u],v)-\omega(u\bullet v,w).$
@@
Eqns. (2) and (3) imply that the multiplication operator $\mathrm{L}^{\bullet}_{u}$, is $\omega$-symmetric, for every $u\in\mathfrak{g}$, and that $\mathop{\mathrm{ad}}\nolimits_{u}=\mathrm{L}^{\bullet}_{u}-\mathrm{R}^{\bullet}_{u}$, where $\mathop{\mathrm{ad}}\nolimits_{u}:\mathfrak{g}\to\mathfrak{g}$ is the adjoint operator. Let us denote by $Z(\mathfrak{g})$ the center of $\mathfrak{g}$}. The following relations hold (where $\mathfrak{g}\bullet\mathfrak{g}={u\bullet v\mid u,v\in\mathfrak{g}}$):
@@
Following [BBE, BE, BM], if a pre-symplectic Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ satisfies the condition (6), then it is called a quasi-Frobenius Lie superalgebra. Cohomologically speaking, this condition means that $\omega\in Z^{2}({\mathfrak{g}};\operatorname{\mathbb{K}})$. In the case where $\omega$ is a coboundary, we call $({\mathfrak{g}},[\;.\;],\omega)$ a Frobenius Lie superalgebra, see [E, O]. Furthermore, if $\omega$ is orthosymplectic (resp. periplectic), we call $(\mathfrak{g},[\;,\;],\omega)$ an orthosymplectic (resp. periplectic) quasi-Frobenuis Lie superalgebra.
@@ 3. $(iii)$
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius nilpotent Lie superalgebra. Then $\mathfrak{g},\mathfrak{g}]$ and $Z(\mathfrak{g})$ are both degenerate.
@@
$u\circ v=-(-1)^{|u||v|}v\bullet u,$
for all $u,v\in\mathfrak{g}$. It is clear that $(\mathfrak{g},\circ)$ is a left-symmetric superalgebra. Moreover, Prop. 3.1.5.  Proposition implies that
(8) $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),$
for all $u,v,w\in\mathfrak{g}$.
@@ 1. $(i)$
    (9) $\omega({\mathrm{R}}_{v}^{\circ}(u),w)=(-1)^{|v||w|}\omega(u,{\mathrm{R}}_{v}^{\circ}(w)),\quad\forall u,v,w\in{\mathcal{A}}.$
@@ 2. $(ii)$
    $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\quad\forall u,v,w\in{\mathcal{A}}.$
@@ 3.2. The symplectic product
$\begin{array}[]{lcrlcrlcrlcrlcrlcr}e_{1}\star e_{1}&=&-\tfrac{1}{3}e_{1},&e_{1}\star e_{2}&=&-\tfrac{1}{3}e_{2},&e_{1}\star f_{1}&=&-\tfrac{1}{3}f_{1},&e_{1}\star f_{2}&=&-\tfrac{1}{3}f_{2},\\[3.0pt] e_{2}\star e_{1}&=&\tfrac{1}{3}e_{2},&e_{2}\star f_{2}&=&\tfrac{1}{3}f_{1},&f_{1}\star e_{1}&=&-\tfrac{2}{3}f_{1},&f_{2}\star e_{1}&=&-\tfrac{2}{3}f_{2},\\[3.0pt] f_{2}\star e_{2}&=&-\tfrac{2}{3}f_{2}.\end{array}$
@@
$e_{1},f_{1}]=f_{2},\qquad[f_{1},f_{1}]=e_{1},\quad\text{ and }\quad\omega=2e_{1}^{*}\wedge e_{2}^{*}-f_{1}^{*}\wedge f_{2}^{*}.$
@@
Let $({\mathcal{A}},\star,\omega)$ be a Lie-admissible superalgebra endowed with a non-degenerate bilinear anti-symmetric form such that ${\mathrm{L}}^{\star}_{u}$ is $\omega$-anti-symmetric, for any $u\in{\mathcal{A}}$. Then $({\mathcal{A}}^{-},[\,,\,]_{\star},\omega)$ is a quasi-Frobenius Lie superalgebra.
@@
$\quad-\frac{1}{3}(-1)^{|u||v|}\omega(M(v,u),w))-\frac{1}{3}(-1)^{|u||v|}\omega(v,M(u,w))-\frac{1}{3}(-1)^{|u||v||u||w|}\omega(v,M(w,u)))$
@@
Let $E$ be a vector superspace equipped with an orthosymplectic (resp. periplectic) form $\omega$. Then there exists an even (resp. odd) symmetric bilinear map $f:E\times E\to E$ such that (for all $u,v,w\in E$)
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)).$
@@
$u\bar{\star}v=u\star v+f(u,v),\quad\forall u,v\in\mathfrak{g}$
@@
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)),$
for any $u,v,w\in\mathfrak{g}.$
@@
${}_{\bar{\star}}$ $=u\star v+f(u,v)-(-1)^{|u||v|}v\star u-(-1)^{|u||v|}f(u,v)$
@@
Thus, $\omega(u\bar{\star}v,w)+(-1)^{|u||v|}\omega(v,u\bar{\star}w)=0$ if and only if $\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w))$. ∎
@@
According to Prop. 3.2.5.  Theorem, there exists a symplectic product $\star$. Moreover, by Lemma 3.2.6.  Lemma, there exists an even symmetric bilinear map $f:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$ satisfying the condition
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)).$
Thus, by Prop. 3.2.7.  Proposition, the product defined by
$u\bar{\star}v=u\star v+f(u,v),\quad\forall u,v\in\mathfrak{g},$
@@
In [BB], the authors introduced a symplectic connection that depends on the Lie bracket and the symplectic form. In this work, we introduce a natural symplectic product for quasi-Frobenius Lie superalgebras that depends on both the Lie bracket and the form, and we establish its uniqueness in Prop. 3.2.9.  Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra. Following the same procedure as in Prop. 3.2.5.  Theorem, and using Eq. (8), we define a product $\circ$ on $\mathfrak{g}$ by
$\omega(\mathrm{L}_{u}^{\circ}(v),w)=(-1)^{|u||v|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\qquad\forall u,v,w\in\mathfrak{g}.$
@@
Since the form $\omega$ is closed, this simplifies to $\omega([u,v],w)=3a\omega([u,v],w),$ and thus $a=\frac{1}{3}$. Therefore, the unique symplectic product compatible with the bracket [,][\;,\;] and the form $\omega$ is (for all $u,v,w\in\mathfrak{g}$):
$\omega(u\star v,w)=\frac{1}{3}\left(\omega([u,v],w)-(-1)^{|v||w|}\omega([u,w],v)\right).\qed$

### v2

# arXiv:2511.20230v2[math.RA] 28 Feb 2026

@@ 1.1. The Riemannian setting
The curvature of $(\mathfrak{g},[\;,\;],\langle\;,\;\rangle)$ at the identity is
@@
A pseudo-Riemannian Lie group $(G,\mu)$ is called flat when its curvature vanishes identically; in this case, $(\mathfrak{g},[\;,\;],\langle\;,\;\rangle)$ will be referred to as a flat pseudo-Riemannian Lie algebra. In fact, flatness is equivalent to $(\mathfrak{g},\star)$ being a left-symmetric algebra.
@@
As in the pseudo-Riemannian case, one defines the curvature of this connection. A symplectic Lie group $(G,\Omega)$ is called flat when its curvature vanishes identically. In this case, the induced Lie algebra $(\mathfrak{g},[\;,\;],\omega)$ is a flat quasi-Frobenius Lie algebra. In fact, flatness is equivalent to the fact that the symplectic product defines a left-symmetric structure on $\mathfrak{g}$.
@@ 1.3. Outline of the paper
for all homogeneous $u,v,w\in{\mathcal{A}}$. The product $\bullet$ is also characterized by a Koszul formula (see Eq. (4)). Its existence and uniqueness are established in Prop 3.1.1.  Proposition.
Further, we study the Levi-Civita product in the case where the pre-symplectic superalgebra is a Lie superalgebra and the form $\omega$ is closed. We call these superalgebras quasi-Frobenius Lie superalgebras, following [BBE, BE, BM]. It is worth mentioning that several authors call them symplectic Lie superalgebras (see [BC, BEB, BBNS] to cite a few). We show that the Levi-Civita product associated with a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is always right-symmetric (see Prop. 3.1.6.  Proposition) and satisfies
@@
In addition, we introduce a second product $\star$ on a quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$, defined by
@@
We show nevertheless that in any quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ we can construct a unique symplectic product with the key property that it depends only on the Lie bracket and the symplectic form $\omega$. This product is explicitly given by:
@@
In Section 6, we show that all non-abelian flat quasi-Frobenius Lie superalgebras have a degenerate center, see Theorem 6.1.  Theorem. Moreover, every such Lie superalgebra can be obtained via flat double extensions. In particular, every flat orthosymplectic quasi-Frobenius Lie superalgebra can be constructed by a sequence of flat double extensions starting from an abelian Lie superalgebra (see Corollary 6.8.  Corollary), while flat periplectic quasi-Frobenius Lie superalgebras can be obtained by a sequence of flat double extensions starting from the trivial algebra $\{0\}$ (see Corollary 6.9.2.  Corollary). Furthermore, we show that every flat orthosymplectic quasi-Frobenius Lie superalgebra is not only solvable but nilpotent (Theorem 6.9.  Theorem).
@@ 2. Backgrounds
$\omega(u,v)=-(-1)^{|u||v|}\omega(v,u),\text{ for all homogeneous }u,v\in V.$
@@
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a finite-dimensional superspace with $V_{\bar{1}}\neq\{0\}$, endowed with a periplectic form $\omega$. Then $\dim(V_{\bar{0}})=\dim(V_{\bar{1}})$.
@@
$N_{\ell}({\mathcal{A}},\bullet):=\left\{u\in{\mathcal{A}},\mathrm{L}^{\bullet}_{u}=0\right\},\;N_{r}({\mathcal{A}},\bullet):=\left\{u\in{\mathcal{A}},\mathrm{R}_{u}^{\bullet}=0\right\},\;\text{and}\;N({\mathcal{A}},\bullet):=N_{\ell}({\mathcal{A}},\bullet)\cap N_{r}({\mathcal{A}},\bullet).$
@@
A superalgebra $(\mathcal{A},[\;,\;])$ is said to be a Lie superalgebra if the following conditions hold for all homogeneous elements $x,y,z\in{\mathcal{A}}:$
  • (Super anti-commutativity):
    $x,y]=-(-1)^{|x||y|}[y,x].$
  • (Super Jacobi identity):
    $x,y],z]=[x,[y,z]]-(-1)^{|x||y|}\,[y,[x,z]].$
We recall that the super Jacobi identity generalizes the classical Jacobi identity by incorporating the parity signs arising from the ${\mathbb{Z}}_{2}$-grading. In particular, when all elements are even, it reduces to the usual Jacobi identity. The identity ensures that the adjoint action defines a representation of the Lie superalgebra.
Throughout this paper, a general superalgebra will be denoted by $({\mathcal{A}},\bullet)$, whereas a Lie superalgebra will be denoted by $(\mathfrak{g},[\;,\;])$. For a thorough exposition of Lie superalgebras, we refer the reader to [L, S].
@@
Example.
Consider the superalgebra ${\mathcal{A}}_{n}$ defined on the homogeneous basis $(e_{1},\dots,e_{n}\mid f_{1},\dots,f_{n}),$ whose non-zero products are given by
$e_{1}\bullet e_{i}$ $=e_{i+1},$ $e_{1}\bullet f_{j}$ $=f_{j+1},$ $2\leq i\leq n-1,\,1\leq j\leq n-1,$
$e_{1}\bullet e_{1}$ $=(-1)^{n+1}e_{2},$ $e_{i}\bullet f_{\,n-i}$ $=f_{\,n-i}\bullet e_{i}=(-1)^{n-i}\,f_{n},$ $2\leq i\leq n-1,$
A straightforward computation shows that $({\mathcal{A}}_{n},\bullet)$ is a left-symmetric superalgebra.
The sub-adjacent Lie superalgebra ${\mathcal{A}}_{n}^{-}$ is given by
$[e_{1},e_{i}]=e_{i+1},\quad[e_{1},f_{i}]=f_{i+1},\quad 1\leq i\leq n-1.$
@@ 3.1. The Levi-Civita product
(Uniqueness) Suppose that there exist two products $\bullet$ and $\bullet^{\prime}$ satisfying conditions (2), (3). For any $u,v,w\in\mathfrak{g}$, we have
$\begin{array}[]{lcl}\omega(u\bullet v,w)&=&\omega([u,v],w)+(-1)^{|u||v|}\omega(v\bullet u,w)\\[5.69054pt] &=&\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega(v\bullet w,u)\\[5.69054pt] &=&\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)-(-1)^{|u||v|+|w||u|+|v||w|}\omega(w\bullet v,u)\\[5.69054pt] &=&\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega(w\bullet u,v)\\[5.69054pt] &=&\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega([w,u],v)\\[5.69054pt] &&-\omega(u\bullet v,w).\end{array}$
@@
Eqns. (2) and (3) imply that the multiplication operator $\mathrm{L}^{\bullet}_{u}$, is $\omega$-symmetric, for every $u\in\mathfrak{g}$, and that $\mathop{\mathrm{ad}}\nolimits_{u}=\mathrm{L}^{\bullet}_{u}-\mathrm{R}^{\bullet}_{u}$, where $\mathop{\mathrm{ad}}\nolimits_{u}:\mathfrak{g}\to\mathfrak{g}$ is the adjoint operator. Let us denote by $Z(\mathfrak{g})$ the center of $\mathfrak{g}$}. The following relations hold (where $\mathfrak{g}\bullet\mathfrak{g}=\{u\bullet v\mid u,v\in\mathfrak{g}\}$):
@@
Following [BBE, BE, BM], if a pre-symplectic Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ satisfies the condition (6), then it is called a quasi-Frobenius Lie superalgebra. Cohomologically speaking, this condition means that $\omega\in Z^{2}({\mathfrak{g}};\operatorname{\mathbb{K}})$. In the case where $\omega$ is a coboundary, we call $({\mathfrak{g}},[\;,\;],\omega)$ a Frobenius Lie superalgebra, see [E, O]. Furthermore, if $\omega$ is orthosymplectic (resp. periplectic), we call $(\mathfrak{g},[\;,\;],\omega)$ an orthosymplectic (resp. periplectic) quasi-Frobenuis Lie superalgebra.
@@ 3. $(iii)$
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius nilpotent Lie superalgebra. Then, $\mathfrak{g},\mathfrak{g}]$ and $Z(\mathfrak{g})$ are both degenerate.
@@
$u\circ v=-(-1)^{|u||v|}v\bullet u,\quad\text{for all $u,v\in\mathfrak{g}$.}$
It is clear that $(\mathfrak{g},\circ)$ is a left-symmetric superalgebra. Moreover, Prop. 3.1.6.  Proposition implies that
(8) $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\quad\text{for all $u,v,w\in\mathfrak{g}$.}$
@@ 1. $(i)$
    (9) $\omega({\mathrm{R}}_{v}^{\circ}(u),w)=(-1)^{|v||w|}\omega(u,{\mathrm{R}}_{v}^{\circ}(w)),\quad\text{for all }u,v,w\in{\mathcal{A}}.$
@@ 2. $(ii)$
    $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\quad\text{for all }u,v,w\in{\mathcal{A}}.$
@@ 3.2. The symplectic product
$\begin{array}[]{lcrlcrlcrlcrlcrlcr}e_{1}\star e_{1}&=&-\tfrac{1}{3}e_{1},&e_{1}\star e_{2}&=&-\tfrac{1}{3}e_{2},&e_{1}\star f_{1}&=&\tfrac{1}{3}f_{1},&e_{1}\star f_{2}&=&\tfrac{1}{3}f_{2},\\[3.0pt] e_{2}\star e_{1}&=&-\tfrac{1}{3}e_{2},&e_{2}\star f_{2}&=&\tfrac{1}{3}f_{1},&f_{1}\star e_{1}&=&-\tfrac{2}{3}f_{1},&f_{2}\star e_{1}&=&-\tfrac{2}{3}f_{2},\\[3.0pt] f_{2}\star e_{2}&=&-\tfrac{2}{3}f_{1}.\end{array}$
@@
$e_{1},f_{1}]=f_{2},\qquad[f_{1},f_{1}]=e_{2},\quad\text{ and }\quad\omega=2e_{1}^{*}\wedge e_{2}^{*}-f_{1}^{*}\wedge f_{2}^{*}.$
@@
Remark.
We adopt the notations $D^{5}$ and $C_{3}+A$ from Backhouse [B], who classified all real Lie superalgebras of dimension 4. The reader is referred to [B] for an explanation of this nomenclature.
@@
Let $({\mathcal{A}},\star)$ be a Lie-admissible superalgebra equipped with a non-degenerate anti-symmetric bilinear form $\omega$ such that ${\mathrm{L}}^{\star}_{u}$ is $\omega$-anti-symmetric, for any $u\in{\mathcal{A}}$. Then the Lie superalgebra $({\mathcal{A}}^{-},[\;,\;]_{\star},\omega)$ is quasi-Frobenius.
@@
$\quad-\frac{1}{3}(-1)^{|u||v|}\omega(M(v,u),w)-\frac{1}{3}(-1)^{|u||v|}\omega(v,M(u,w))-\frac{1}{3}(-1)^{|u||v||u||w|}\omega(v,M(w,u))$
@@
Let $E$ be a vector superspace equipped with an orthosymplectic (resp. periplectic) form $\omega$. Then there exists an even (resp. odd) symmetric bilinear map $f:E\times E\to E$ such that
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)),\quad\textup{for all $u,v,w\in E$}.$
@@
$u\bar{\star}v=u\star v+f(u,v),\quad\text{for all }u,v\in\mathfrak{g}$
@@
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)),\;\text{for all $u,v,w\in\mathfrak{g}.$}$
@@
$[u,v]_{\bar{\star}}$ $=u\star v+f(u,v)-(-1)^{|u||v|}v\star u-(-1)^{|u||v|}f(u,v)$
@@
Thus,
$\omega(u\bar{\star}v,w)+(-1)^{|u||v|}\omega(v,u\bar{\star}w)=0\text{ if and only if }\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)).$
@@
According to Prop. 3.2.7.  Theorem, there exists a symplectic product $\star$. Moreover, by Lemma 3.2.8.  Lemma, there exists an even symmetric bilinear map $f:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$ satisfying the condition
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)),\quad\text{ for all }u,v,w\in\mathfrak{g}.$
Thus, by Prop. 3.2.9.  Proposition, the product defined by
$u\bar{\star}v=u\star v+f(u,v),\quad\text{ for all }u,v\in\mathfrak{g},$
@@
In [BB], the authors introduced a symplectic connection that depends on the Lie bracket and the symplectic form. In this work, we introduce a natural symplectic product for quasi-Frobenius Lie superalgebras that depends on both the Lie bracket and the form, and we establish its uniqueness in Prop. 3.2.11.  Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra. Following the same procedure as in Prop. 3.2.7.  Theorem, and using Eq. (8), we define a product $\circ$ on $\mathfrak{g}$ by
$\omega(\mathrm{L}_{u}^{\circ}(v),w)=(-1)^{|u||v|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\qquad\text{for all }u,v,w\in\mathfrak{g}.$
@@
Since the form $\omega$ is closed, this simplifies to $\omega([u,v],w)=3a\omega([u,v],w),$ and thus $a=\frac{1}{3}$. Therefore, the unique symplectic product compatible with the bracket [,][\;,\;] and the form $\omega$ is:
$\omega(u\star v,w)=\frac{1}{3}\left(\omega([u,v],w)+(-1)^{|v||w|}\omega([u,w],v)\right),\quad\text{for all $u,v,w\in\mathfrak{g}$.}\qed$